### 13 Jan 2024 v1

# Bridging the Preference Gap between Retrievers and LLMs

Abstract
Large Language Models (LLMs) have demonstrated superior results across a wide range of tasks, while retrieval has long been established as an effective means of obtaining task-relevant information for humans. Retrieval-augmented Generation (RAG) are known for their effectiveness in knowledge-intensive tasks by locating relevant information and placing it within the context window of the LLM. However, the relationship between retrievers and LLMs is still under-investigated. Most existing work treats the retriever and the LLM as independent components and leaves a gap between retrieving human-friendly information and assembling a LLM-friendly context. In this work, we examine a novel bridge model, validate the ranking and selection assumptions in retrievers in the context of RAG, and propose a training framework that chains together supervised and reinforcement learning to learn a bridge model. Empirical results demonstrate the effectiveness of our method in both question-answering and personalized generation tasks.
Bridging the Preference Gap between Retrievers and LLMs
[…]
1 Introduction
Large language models (LLMs) such as GPT-3 Brown et al. (2020) and PaLM2 Anil et al. (2023), have demonstrated impressive performance on a wide range of language tasks. Alongside LLMs, retrieval-augmented generation (RAG), which can retrieve knowledge from an external dataset when needed, have also produced strong results in many knowledge intensive-tasks.
Figure 1: We observe a preference gap when changing the ranking and selection within the LLM’s input context. The experiments are conducted with GTR Ni et al. (2021) as retriever and a fixed Palm2-S as the LLMs. “Top-K” means we select the top-k GTR retrieved passages. Different colors indicate different datasets. The selected passages are used as additional context. We see that selection (Top-1) significantly affects (either positively or negatively) the model’s performance, while ranking (randomized top-5) does not have much impact (the metrics are detailed in Sec. 5.2).
However, despite some extensive RAG studies Khandelwal et al. (2020); Borgeaud et al. (2022); Izacard et al. (2022); Yasunaga et al. (2023), most of them study retrievers and LLMs separately. On one hand, considerable effort has been dedicated to designing user-friendly retrievers, based on the general belief that ranking is paramount, as humans typically read from top to bottom. On the other hand, LLMs exhibit preferences different from humans and yield accurate predictions only when prompts align with these preferences. This discrepancy leads to sub-optimal design in current RAG systems, a phenomenon we term preference gap. This gap manifests in various aspects. For example, the general belief on ranking may not align with LLM’s preferences due to the self-attention mechanism in Transformers, which can focus on any token regardless of its sequential position. Another aspect is selection; humans can disregard irrelevant passages, but it has been shown that LLMs are highly sensitive to irrelevant content Shi et al. (2023a). There are likely more gaps that further diverge from human behavior, such as repetition, which is generally considered detrimental for retrieval systems Xia et al. (2017), but maybe be useful to LLMs for relevance weighting.
We empirically investigate this preference gap, focusing specifically on ranking and selection. As shown in Fig. 1, the performance varied by around 1% when we randomized the top-5 retrieved passages [^1] We note that this is different from Liu et al. (2023), where a “loss in the middle” phenomenon is observed, meaning that the model are better at using the relevant information at the beginning or end of its input context. This is probably due to different number of passages used. In Liu et al. (2023), 20 total documents are employed, whereas we used 5 passages. Regardless, this does not affect the conclusion that there exists a preference gap between the retrievers and LLMs which needs to be bridged.. However, the variation exceeded 5% when selecting the top-1 passage. This indicates the general belief in the importance of ranking in retrieval does not apply to LLMs , indicating that selection could be more crucial. This finding confirms the existence of the preference gap between the retrievers and LLMs, and highlights the importance of bridging this preference gap to enhance RAG. This is a crucial insight that has not been revealed before, which can guide RAG designers towards what should be achieved in the model.
To design a RAG model, existing work has tried to finetune the LLMs to align with the retriever’s preference, or to finetune the retriever to align with the LLM’s preference. However, since they primarily focuses on ranking, their approaches are sub-optimal for bridging the preference gap. Fine-tuning retrievers can only achieve re-ranking, and fine-tuning LLMs is often impractical, as many are only accessible through APIs. To better bridge the gap beyond ranking, we propose a framework called BGM (Bridging the Gap between retrievers and LLMs). We propose to keep the retriever and LLM fixed, and train a bridge model in between. The bridge model aims to transform the retrieved information into a format that LLMs can work with effectively.
Since the preference gap is present in both ranking and selection, and potentially extends to aspects such as repetition, in contrast to the conventional scoring format, we structure the bridge model as a sequence-to-sequence (seq2seq) model. This approach allows it to not only rerank but also dynamically select passages for each query, and potentially employ more advanced strategies like repetition. In a typical RAG, there is no ground truth relevance label for what should be retrieved but only ground truth label for the downstream task. Existing supervised learning (SL) approaches use the supervision provided by the LLM, such as the perplexity of downstream tasks. However, this is ineffective due to sparse supervision as it is nearly impossible to feed all possible retrieved sequences into the LLM to obtain supervision. Even worse, such SL relies on the intermediate relevance label and is not end-to-end training on the downstream tasks. To address these issues, we employed reinforcement learning (RL) on the SL trained bridge model, where the downstream performance metrics are used as the reward and the bridge model are regarded as a policy model. This framework, chaining SL and RL, provides increased supervision from the downstream task. It also offers the model the flexibility to explore more advanced strategies, such as repetition, in forming the optimal passage sequence.
Our experiments reveal that BGM can enhance the performance of various downstream tasks, such as Question Answering (QA) and personalized generation, across a spectrum of datasets, from public QA and amazon reviews to private email conversations. Notably, the modified passages retrieved by BGM surpass the performance of strong retrievers and reranking baseline models. This underscores the significance and promise of the bridging approach in the realm of RAG. In summary, our contributions can be summarized as follows:
  • We empirically establish the existence of the preference gap between retrievers and LLMs, and introduce BGM, which designed to address this preference gap.
  • We propose a seq2seq bridge model to jointly accomplish reranking and selection, adapting the retrieved information to be LLM-friendly. We employ a SL and RL training scheme to optimize this adaptation process.
  • The evaluation of BGM comprises diverse datasets, including QA and text generation, including both publicly available and personal information sources. The comprehensive evaluation underscores the effectiveness of BGM in bridging the preference gap and improving performance in downstream tasks.
Figure 2: Differing from previous RAG that update either LLMs, retrievers, or both, BGM not only treats the LLM as a black box but also freezes the retrieval. Only a lightweight bridge model (a Seq2seq LM) is tunable in order to adapt the retrieved information to satisfy the LLM’s preference. This makes BGM applicable to large LLMs, which are often served via APIs and also avoids expensive updates to the retriever.
2 Related Work
Retrieval-augmented Generation (RAG). Augmenting LLMs with relevant information retrieved from various knowledge sources has shown effective in improving performance across numerous NLP tasks, including language modeling Borgeaud et al. (2022); Khandelwal et al. (2020); Shi et al. (2023b), question answering Lewis et al. (2020); Izacard et al. (2022); de Jong et al. (2023); De Jong et al. (2023); Shi et al. (2023b); Guu et al. (2020); Izacard and Grave (2020); Xu et al. (2023), fact versification Lewis et al. (2020) and text generation Lewis et al. (2020). Specifically, RAG utilizes input as a query and comprises two main components: (1) retriever retrieves a set of items from a large corpus. Note that the retrieval units vary across different works, including documents, passages, or even tokens. In this study, we focus on retrieving passages; and (2) LLM incorporate the retrieved passages as additional information, integrating them into the input context for making final predictions.
A fundamental question in this process arises regarding the differing preferences between LLMs and existing retrievers, with LLMs performing optimally only when their preferences are satisfied. Bridging the gap between these differing preferences is crucial. Depending on which components are subject to updates, this challenge can be categorized into three families
Finetuning retrievers and LLMs jointly. This is the most widely used and conventional setting of RAG Izacard et al. (2022); Khandelwal et al. (2020); Wu et al. (2022); Guu et al. (2020); Lewis et al. (2020). However, they are mostly less than 1B so may nor regard as “large” LM. For example, Altas Izacard et al. (2022) finetunes LLM (T5 Raffel et al. (2020a)) and retriever (Contriever Izacard et al. (2021)) jointly by leveraging the LLM to provide supervisory signal to train the retriever. RAG Lewis et al. (2020) uses a tunnable query encoder and DPR Karpukhin et al. (2020) as retriever, BART as LLM, and design an end-to-end training and inference schema to train the query encoder and the LLM.
Finetuning LLMs only. Updating retrievers is not always desirable as it is costly and requires the document index to be periodically updated. To bridge the preference gap, it is also possible to only update the LLMs. Fid Izacard and Grave (2020) takes the retrieved documents and query as input, finetunes the LLM to adapt to the external information. Similarly, Lummen De Jong et al. (2023) and Glimmer de Jong et al. (2023) improve FiD via adding reranker and pre-encoding memeory.
Finetuning retrievers only. Although above systems have shown improves results, they are not always applicable in practice. Many LLMs such as GPT-3 and Codex are not open-sourced due to commercial considerations and are only available as black-box APIs. Through these APIs, users can send queries and receive responses. Given this constraint, a natural approach is to only update the retrievers to ensure they can retrieve passages that are more compatible with LLMs. REPLUG Shi et al. (2023b) adapts a similar idea as Atlas but fix the LM. RECOMP Xu et al. (2023) trains a compressors to summarize the retrieved document from retriever. However, this family of models is incapable of performing any sample-level selection and can only choose top passages by setting a fixed threshold.
Unlike the existing three families, BGM works on LLM (> 11B) and directly working on bridging the preference gap, without finetuning LLMs or retrievers but a bridge model in between (Fig. 2).
RL for information retrieval. Before the LLM era, RL has been used in information retrieval (IR) Xia et al. (2017); Wei et al. (2017); Zeng et al. (2018); Xu et al. (2020). The core approach was to frame the IR problem as a Markov Decision Process and apply an RL algorithm to solve it. Typically, an IR task would be structured to determine which document to select for a ranking position, using ranking metrics such as DCG as the reward. This approach had advantages, including using the ranking position directly as supervision and optimizing the IR measure directly. However, none of these existing works explored the application of RL in the context of RAG.
In the LLM era, RL has been used in query rewriting for retrieval Wu et al. (2021); Nogueira and Cho (2017); Adolphs et al. (2021), where a black-box retriever is assumed. This is a different problem from RAG. Bacciu et al. (2023) suggest using RL to fine-tune retriever in the RAG context in their opinion paper, but it does not conduct any experiments. More crucially, their work fails to recognize the importance of bridging the gap between retrievers and LLMs, nor does it specify what the bridge model should be.
3 Problem Formulation
Retriever. Given an input $x$, the retriever aims to retrieve a ranked list of passages from a corpus $D=\{d_{i}\}_{i=1}^{m}$ that are relevant to $x$. In this work, we assume a frozen dense retriever. Typically, a dual encoder architecture is applied, where an encoder is used to encode both the input context $x$ and the passage $d$. Specifically, the encoder maps each passage to an embedding $\bm{E}(d)$. The similarity between input and passage embedding is computed by their cosine similarity,
$s(d,x)=\text{cos}(\bm{E}(d),\bm{E}(x)).$ (1)
The top-k passages that have the highest similarity scores when compared with the input $x$ are retrieved in this step.
$(d^{\text{retr.}}_{j})_{j=1}^{k}=\text{Top-K}(\{s(d,x)\}_{i=1}^{m}).$ (2)
Bridge Model for RAG. The retrieved top-K passages provide richer information about the original input $x$ and can potentially help the LLM to make a better prediction on downstream tasks. However, since there is preference gap, we propose a bridge model $\bm{B}$ to adapt the retrieved passages to LLM-friendly passages. As mentioned in the Sec. 1, the bridge model is a seq2seq model. It takes as input all the retrieved passages $(d^{\text{retr.}}_{j})_{j=1}^{k}$ and output a subset of adapted passages $(d^{\text{bdr.}}_{j})_{j=1}^{n}$. This format is advantageous because the seq2seq model automatically considers ranking by generating the next token based on the preceding one and selection by placing the end-of-sentence token in the appropriate position. Note that $n$ may or may not equal to $k$ due to selection.
$(d^{\text{bdr.}}_{j})_{j=1}^{n}=\bm{B}((d^{\text{retr.}}_{j})_{j=1}^{k}).$ (3)
It is important to note that in practice, the bridge model’s input includes the query, passage IDs, and passage content, while the output consists of passage IDs. We then convert the obtained passage IDs to their corresponding documents when we use them in RL (Sec. 4.2). Additionally, the passages in passage sequence (e.g., $(d^{\text{retr.}}_{j})_{j=1}^{n}$) are concatenated into a long sequence before being fed to the network.
Retrieval-augmented Generation with bridge model. Given the adapted passages from the bridge model as context, $(d^{\text{bdr.}}_{j})_{j=1}^{n}$, we concatenate them with the input $x$, and fed the resulting long sequence into the LLM to obtain the output for downstream tasks.
Figure 3: Illustrating the training of BGM: Step 1: First, we prepare the silver passage sequence for supervised learning (SL) through a greedy search on the retrieved passages. Step 2: This silver passage sequence is then used for the supervised training of the bridge model. The bridge model is a seq2seq model that takes the query, passage ID, and passage content as input and outputs the passage IDs. Step 3: Finally, the trained bridge model is treated as a policy model and is further trained using reinforcement learning
4 Training the Bridge Model
In Eq. 3, we format the bridge model as seq2seq model. However, it is challenging to effectively train it as there is no ground truth retrieved passages and LLM is assumed to be API without gradient. To address the issue, we propose to chain supervised learning (SL) and reinforcement learning (RL), where SL aims to reduce the search space of RL and provides a reasonably good initial model that does ranking and selection. Meanwhile, RL aims to optimize the policy model, i.e., bridge model, against the reward.
4.1 Supervised Learning
To conduct SL, the ground-truth sequence is required for each query. In a typical RAG, the ground truth relevance passage for the query is not provided, let alone the ground-truth sequence. To address this, we propose to synthesis silver passage sequence by selecting only the useful passage. This is done by greedy search that incrementally selects the next passage that can maximize the downstream task performance.
Synthesising silver passage sequence using greedy search. Assume that the downstream task performance for a given passage sequence of any length, used as context, is denoted as $R(\cdot)$. The empty passage sequence, i.e., no retrieved passage, denoted as $d^{\text{retr.}}_{\varnothing}$, has downstream task performance $R(d^{\text{retr.}}_{\varnothing})$. Starting from $d^{\text{retr.}}_{\varnothing}$ and empty silver passage sequence $d^{\text{silv.}}$, we try to add (i.e., concatenate) the candidate passage that can improve most to the silver passage sequence, until no improvement can be made. Algorithm 1 shows the pseudo-code for synthesising silver passage sequence.
Input: $(d^{\text{bdr.}}_{j})_{j=1}^{n}$, $R(\cdot)$
Output: $(d_{j}^{\text{silv.}})_{j=1}^{s}$
1 $d^{\text{silv.}}=()$;
2 $silv=R(d^{\text{retr.}}_{\varnothing})$;
3 $prev=0$;
4 $best=\varnothing$;
5 while True do
6       for $d\leftarrow d^{\text{retr.}}_{1}$ to $d^{\text{retr.}}_{n}$ where $d$ not in $d^{\text{silv.}}$ do
7             $seq$ = concat($d$, $d^{\text{silv.}}$);
8             $cur$ = $R(seq)$;
9             if (cur > silv) and (cur > prev) then
10                   $best$ = $d$;
11                   $prev$ = $cur$;
12
13
14      if best != $\varnothing$ then
15             add $best$ to the end of $d^{\text{silv.}}$;
16             $silv=prev$;
17             $prev=0$;
18             $best=\varnothing$
19      else
20             Break;
21
22
Algorithm 1 Synthesis Silver Passage Sequence
Training. Training with the silver passage sequence is straightforward, achieved by applying cross-entropy loss.
4.2 Reinforcement Learning
Although supervised learning can already help training the bridge model, it is still ineffective (we can see using SL alone results in mixed performance compared to using a retriever alone in Table 5). This is attributed to sparse supervision and the lack of end-to-end training on downstream results.
To address these issues, we apply RL to further train the bridge model. RL does not limit the possible passage sequences (we only consider permutations or deletions in the silver passage sequence in SL. However, the optimal passage sequence might require more complex manipulations, such as repetition, which RL can accommodate) and give more supervision to find the best passages sequence. Using the performance of downstream task as the reward, the bridge model is trained in an end-to-end manner on the downstream task. Specifically, our task can be formulated as a RL problem as follows:
Reward is the performance of tasks against their ground-truth labels as we focus on labeled downstream tasks.
Policy Model is the bridge model that need to be trained.
Action Space is limitted to Passage IDs as we are interested in adapting the retrieved passages
Training. The reward objective can be optimized by any off-the-shelf RL algorithm, e.g., proximal policy optimization (PPO).
5 Experiments
5.1 Datasets and Baselines
Datasets. We consider four datasets, ranging from popular QA datasets to less investigated personalized ones. We also include one dataset that contains private email conversations (Avocado Email), which is unlikely to be included in the LLM’s pre-training datasets. This will further help us investigate the effectiveness of our proposed BGM model, as the LLM will have to rely on the retrieved passages. The summary of statistics is given in Table 1.
Open-domain QA. We conduct evaluations on two open-domain QA: Natural Questions (NQ) Kwiatkowski et al. (2019) and HotpotQA Yang et al. (2018). They both consist of questions, answers collected from Wikipedia and the Web. HotpotQA is a multi-hop QA dataset which requires findings and reasoning over multiple passages to answer the question. The candidate passages are retrieved from WikiPedia pages
Personalized Generation. We follow Li et al. (2023) to construct the personalized generation datasets. Specifically, suppose a user is writing a document, which we call the current document. Given the immediate context $x$ and the user’s personal context, we aim the finish the document as if the user had completed it. The immediate context $x$ or the input to the personalized writing task, is defined as the title and the start of the current document. The candidate passages are retrieved from documents authored by this user in the past. This includes datasets from two domains: Avocado Email (Email) Oard et al. (2015) and Amazon Book (Book) Ni et al. (2019).
Figure 4: An example from the Book dataset. The query consists of the instruction, the document (i.e., review for Book) title, document product (there will be no product information for Email), document start. The target is to complete the text given the query.
Baselines. We consider the state-of-the-art baselines: (1) GTR Ni et al. (2021) a widely recognized retriever that operates independently of LLMs; (2) a variant of GTR, termed Random, in which the order of passages retrieved by GTR is randomized; (3) Point-wise score ranking (PSR) Izacard et al. (2022). This is a variant from Izacard et al. (2022) where we substitute the decoder with our LLM and apply point-wise scoring reranking. This approach can be regarded as utilizing a reranker as a bridge model, but lacking the capability for dynamic selection; and (4) Additionally, we include a non-retrieval baseline, Naive, where no retrieval is conducted to aid in the generation process.
#Training #Val. #Test Avg. # words
NQ 79,168 8,757 3,610 517.82
HotpotQA 68,659 5,600 5,600 564.83
Email 13,305 764 1,227 173.85
Book 20,789 41,331 41,331 124.52
Table 1: Statistic of the 4 datasets.
5.2 LLM and Hyperparamters
We select the T5-XXL (11B) Raffel et al. (2020b) model as our bridge model for most experiments. In the supervised learning stage, the T5-11B model is fine-tuned with a base learning rate of 0.001. A linear warmup scheduler is used for the first 1,000 training steps. Additionally, the square root normalized decay of the learning rate is applied. The model is trained until its performance converges on the validation set. Decoding is performed using beam search with a beam size of 4. We use the PaLM2 model Anil et al. (2023), a new state-of-the-art LLM, as our LLM. It adopts temperature sampling as the decoding strategy. The parameters of PaLM2 are frozen, and we set the temperature to 0 to make the output deterministic. We use Exact-Match (EM) and BLEU as the metric to select the best prompt among prompt variants. It is also used as the reward to train the bridge model in the RL stage. The “Top-K” in Eq. 2 is set to 5.
5.3 Evaluation Results and Analysis
Model NQ HotpotQA Email Book
Metric EM EM BLEU BLEU
Naïve 33.07 28.01 5.57 11.5
Random 43.71 26.10 8.55 8.61
GTR 43.79 25.80 9.76 8.75
PSR 43.60 25.51 9.08 9.14
BGM 45.37 35.64 10.42 12.07
Table 2: Performance of all 4 datasets
Superior of BGM. Table 2 reports the overall performance. We can see that
(1) The proposed BGM outperforms all the 4 baselines in all 4 datasets. This clearly indicates BGM is effective in adapting the retrieved passages.
(2) Compared to the Naive approach, BGM shows significant improvement overall. However, an exception is observed in the Book dataset, where BGM’s improvement is more modest. This suggests that the LLM already possesses a substantial amount of relevant knowledge, likely due to the inclusion of Amazon reviews (which the Book dataset comprises) in the LLM’s pre-training data. Consequently, retrieval is not always essential in this context. This also explains why the Naive approach outperforms other baselines (except BGM) on the Book dataset. It’s important to note that BGM still manages to show improvement, albeit to a lesser extent, in this scenario
(3) Compared to the GTR approach, BGM demonstrates substantial improvement. The NQ dataset shows the least improvement, as most instances only require one retrieved passage, which both GTR and BGM can successfully include. On the other hand, the HotpotQA dataset shows the most significant improvement, indicating that HotpotQA may be more sensitive to irrelevant passages. It’s noteworthy that Random and GTR perform similarly, while GTR shows improvement over Naive (with the exception of the Book dataset). This suggests that the ranking itself does not significantly impact performance, but the selection of passages does
(4) Compared to the PSR approach, BGM once again demonstrates significant improvement. This indicates that pure reranking alone is not sufficient for the bridge model. Selection must also be taken into account. It’s notable that PSR performs similarly to GTR, further suggesting that reranking alone does not greatly affect performance
Model NQ HotpotQA Email Book
Metric EM EM BLEU BLEU
Naive 33.07 28.01 5.57 11.5
PSR 43.6 25.51 9.08 9.14
PSR (Top1) 42.02 32.69 7.28 11.53
PSR (Top2) 42.54 31.05 7.77 10.11
PSR (Top3) 42.85 32.71 8.21 9.70
PSR (Top4) 43.71 32.37 8.26 9.11
BGM 45.37 35.64 10.42 12.07
Table 3: Ablation results - threshold the PSR
Silver Data NQ HotpotQA Email Book
Metric EM EM BLEU BLEU
GTR 43.79 25.8 9.76 8.75
PSR 43.68 29.73 10.1 10.35
Greedy (BGM) 45.37 35.64 10.42 12.07
Table 4: Ablation results - different silver passage sequence for SL
Model NQ HotpotQA Email Book
Metric EM EM BLEU BLEU
GTR 43.79 25.8 9.76 8.75
BGM (SL only) 39.44 34.26 8.62 12.05
BGM 45.37 35.64 10.42 12.07
Table 5: Ablation results - SL only
Model NQ HotpotQA Email Book
Metric EM EM BLEU BLEU
GTR 43.79 25.8 9.76 8.75
FLAN-T5-Large 44.15 35.87 10.18 10.19
FLAN-T5-XL 44.87 35.41 9.64 10.7
FLAN-T5-XXL (BGM) 45.37 35.64 10.42 12.07
Table 6: Ablation results - different bridge model size
Model Palm2-XXS Palm2-S
Data NQ HotpotQA Email Book
Metric EM EM BLEU BLEU
Naive 12.13 14.57 33.07 28.01
Random 31.19 24.41 43.71 26.1
GTR 31.91 23.17 43.79 25.8
PSR 32.04 22.53 43.60 25.51
BGM 39.88 28.69 45.37 35.64
Table 7: Ablation results - different LLM size
Model NQ HotpotQA Email Book
Metric EM EM BLEU BLEU
Test on Palm2-S
BGM 45.37 35.64 10.42 12.07
BGM
(Train on NQ)
33.42 5.66 11.22
BGM
(Train on Email)
35.59 27.98 11.38
Test on Palm2-XXS
BGM
(Train on Palm2-XXS)
39.88 28.69
BGM
(Train on Palm2-S)
30.63 24.55
Table 8: Ablation results - the generability of BGM across various datasets and LLM sizes.
Understanding BGM. We are interested in what happen when we optimize the BGM, whether BGM’s improvement is trival, and how some important setting affects the BGM.
(1) Whether the selection can be achieved naively by manually threshold the PSR? We conducted an ablation experiment using various thresholds for PSR, as shown in Table 3. The term “Top-K” in rows 3 to 6 refers to the selection of only the Top-K passages from PSR reranked passages. The results are mixed, but consistently inferior to those of BGM. This suggests that a naive manual threshold applied to the reranking model is insufficient to meet our objectives. Therefore, it is necessary to consider both reranking and dynamic selection simultaneously
(2) How different silver passage sequence will affect the final performance? In Table 2, we employed greedy search passage sequences as silver passage sequence for SL (Sec. 4.1). However, lacking ground truth passage sequences, we also explored whether this approach is superior to others. Our ablation experiment in Table 4 involved using various types of silver passage sequence. The first row, labeled ’GTR’, indicates the use of GTR’s retrieved passages as silver passage sequence. Similarly, ’PSR’ refers to using PSR reranked passages as silver passage sequence. In the final row, we used greedy search passage sequences, representing the current version of BGM. The results demonstrate that the quality of silver passage sequence significantly affects downstream task performance. While using PSR as silver passage sequence shows improvement over GTR on three datasets, Greedy (BGM) further enhances performance and achieves the best results. Identifying potentially better silver passage sequence is left for future work.
(3) How helpful RL is to the final performance? BGM integrates SL and RL, and we aim to assess the effectiveness of each component. An ablation experiment is detailed in Table 5. We can observe that BGM, when operating with only SL, performs significantly worse than the full BGM model in NQ, HotpotQA, and Email, and in some cases, it even underperforms GTR. This suggests that SL alone is inadequate, highlighting the necessity of incorporating RL. It’s important to note that removing SL is not feasible, as this would result in an excessively large search space for RL to effectively learn.
(4) How different size of bridge model affect the final performance? In Table 2, we utilized Flan-T5-XXL (11B) as the bridge model, which is already smaller than the LLM (24B). However, we are curious about the feasibility of using an even more lightweight language model (LM) as the bridge model. The ablation study shown in Table 6 presents the outcomes of employing bridge models of various sizes. It is evident that all three sizes (large, XL, and XXL) surpass the performance of the setup without a bridge model (i.e., GTR), with the largest size yielding the best results. This demonstrates that a bridge model is beneficial even at a smaller scale, and that larger sizes lead to further improvements.
(5) How different size of LLMs affect the final performance? In Table 2, Palm2-S was used as the LLM. We are interested in evaluating the effectiveness of the bridge model with different sizes of LLMs. We conducted experiments using Palm2-XXS, as detailed in Table 7. It is important to note that the smaller LLM struggles with personalized generation datasets (i.e., Email and Book), resulting in BLEU scores lower than 1%. We opted not to report these results as they may not accurately reflect the trend. However, observations from NQ and HotpotQA suggest that BGM significantly outperforms all baselines by a large margin. This indicates that BGM is effective even with a smaller LLM.
(6) Can a trained bridge model generalize to different datasets and LLMs? The effectiveness of the bridge model across various datasets and LLMs has been demonstrated. A more ambitious objective is to extend this performance to new contexts without additional training. We conducted experiments to investigate the generalizability of the bridge model. The upper section of Table 8 presents the data on generalizability across different datasets. Row 2 shows the performance of BGM (i.e., the bridge model) when trained exclusively on the NQ dataset and then tested on three other unseen datasets. Similarly, row 3 describes training BGM solely on the Email dataset and testing it on the other three. In all cases, the performance is inferior to that of BGM when both training and testing are conducted on the same datasets. This result is expected, as no specific techniques have been developed for dataset generalization, a topic we leave for future work.
In the lower section of Table 8, we present the results of BGM when tested on Palm2-XXS, but trained on Palm2-S. Comparing with results of BGM both trained and tested on Palm2-XXS (as shown in row 1 of the bottom section), the mismatch between LLM sizes leads to a significant decline in performance. This suggests that BGM’s ability to generalize across different LLM sizes is currently limited. Addressing this limitation is considered an important direction for future research.
(8) Case Studies. We provided examples of GTR, PSR, and BGM in Table 9 in the Appendix for the NQ dataset. For question I, both GTR and PSR yield the same incorrect output, even though their contexts have different orders. Only BGM provides the correct answer, indicating that additional irrelevant context can be noisy and detrimental to RAG’s performance. In question II, unlike question I, GTR and PSR produce different answers. This suggests that the ranking can sometimes influence the results. While the specific mechanism by which ranking affects outcomes remains unclear (we regard it as a valuable topic for future research), it is evident that we need to consider ranking, as it may or may not affect performance. Again, BGM selects the most relevant passage and delivers the correct answer. In question III, none of the candidate passages contain the answer (they discuss FaZe Clan and the number of subscribers, but do not identify who has the most subscribers). GTR and PSR provide incorrect answers, as their additional context is unhelpful. In contrast, BGM opts not to select any passages and answers the question using its own parameters, resulting in the correct answer. This demonstrates that retrieval-augmented processes are not always necessary, and BGM is capable of handling such cases
6 Conclusion
This paper highlights the importance of bridging the preference gap between retrievers and LLMs. It begins by demonstrating the existence of this gap, especially in terms of ranking and selection. This provides an important guidance about what should be achieved in a RAG system. It then suggests the development of a lightweight bridge model. This component is designed to adapt user-friendly passages retrieved by the retrievers into formats more suitable for LLMs. To accomplish these objectives, we introduce a novel system named BGM. BGM treats the bridging task as a seq2seq problem, allowing for simultaneous consideration of both ranking and selection. It is trained through a combination of supervised learning and reinforcement learning, facilitating end-to-end training of the bridge model based on the performance in downstream tasks and is capble even when the LLM is black-box (i.e., API). Extensive experiments have demonstrated BGM’s effectiveness.

### 20 Feb 2024 v2

@@
 Abstract
-Large Language Models (LLMs) have demonstrated superior results across a wide range of tasks, while retrieval has long been established as an effective means of obtaining task-relevant information for humans.
-Retrieval-augmented Generation (RAG) are known for their effectiveness in knowledge-intensive tasks by locating relevant information and placing it within the context window of the LLM.
-However, the relationship between retrievers and LLMs is still under-investigated.
-Most existing work treats the retriever and the LLM as independent components and leaves a gap between retrieving human-friendly information and assembling a LLM-friendly context.
-In this work, we examine a novel bridge model, validate the ranking and selection assumptions in retrievers in the context of RAG, and propose a training framework that chains together supervised and reinforcement learning to learn a bridge model.
+Large Language Models (LLMs) have demonstrated superior results across a wide range of tasks, and Retrieval-augmented Generation (RAG) is an effective way to enhance the performance by locating relevant information and placing it into the context window of the LLM.
+However, the relationship between retrievers and LLMs in a RAG is still under-investigated.
+Most existing work treats the retriever and the LLM as independent components and leaves a gap between retrieving human-“friendly” information and assembling a LLM-“friendly” context.
+In this work, we examine a novel bridge mechanism.
+We validate the ranking and selection assumptions of retrievers in the context of RAG and propose a framework that chains together supervised and reinforcement learning to train a bridge model that optimizes the connection between the retriever and the LLM.
 Empirical results demonstrate the effectiveness of our method in both question-answering and personalized generation tasks.
@@
 1 Introduction
-Large language models (LLMs) such as GPT-3 Brown et al.
-(2020) and PaLM2 Anil et al.
-(2023), have demonstrated impressive performance on a wide range of language tasks.
-Alongside LLMs, retrieval-augmented generation (RAG), which can retrieve knowledge from an external dataset when needed, have also produced strong results in many knowledge intensive-tasks.
-Figure 1: We observe a preference gap when changing the ranking and selection within the LLM’s input context.
-The experiments are conducted with GTR Ni et al.
-(2021) as retriever and a fixed Palm2-S as the LLMs.
-“Top-K” means we select the top-k GTR retrieved passages.
-Different colors indicate different datasets.
-The selected passages are used as additional context.
-We see that selection (Top-1) significantly affects (either positively or negatively) the model’s performance, while ranking (randomized top-5) does not have much impact (the metrics are detailed in Sec.
-5.2).
-However, despite some extensive RAG studies Khandelwal et al.
+Large language models (LLMs) such as GPT-4 openai2023gpt4 and PaLM 2 Anil et al.
+(2023), have demonstrated impressive performance on a wide variety of tasks.
+Retrieval-augmented generation (RAG), which retrieves knowledge items from an external data source and puts it into the context window of LLMs, has produced significantly enhanced results in many NLP tasks Khandelwal et al.
 (2020); Borgeaud et al.
@@
 (2022); Yasunaga et al.
-(2023), most of them study retrievers and LLMs separately.
-On one hand, considerable effort has been dedicated to designing user-friendly retrievers, based on the general belief that ranking is paramount, as humans typically read from top to bottom.
-On the other hand, LLMs exhibit preferences different from humans and yield accurate predictions only when prompts align with these preferences.
+(2023).
+Figure 1: We observe a preference gap when alternating the ranking and selection of information in RAG.
+Experiments are conducted with retrieving passages using GTR Ni et al.
+(2021) and using top K of them as additional context for a frozen Palm2-S LLM.
+Different colors indicate different datasets (detailed in Sec.
+5.1) and the Y-axis shows the relative percentage.
+Alternating the selection (Top-1) of information significantly affects (either positively or negatively) the LLM’s performance, while randomizing the ranking of multiple selected items (Top-5) does not have a comparable impact (the metrics are detailed in Sec.
+5.2).
+Note the impact on NQ is even too small to be visible.
+However, most works on RAG study retrievers and LLMs separately.
+On one hand, most retrievers are designed to be human-friendly, usually based on the general belief in classic information retrieval literature that ranking is paramount, as humans typically read from top to bottom robertson1977probability.
+On the other hand, LLMs exhibit preferences different from humans and yield accurate results only when the information in the prompt aligns with these preferences.
 This discrepancy leads to sub-optimal design in current RAG systems, a phenomenon we term preference gap.
 This gap manifests in various aspects.
-For example, the general belief on ranking may not align with LLM’s preferences due to the self-attention mechanism in Transformers, which can focus on any token regardless of its sequential position.
-Another aspect is selection; humans can disregard irrelevant passages, but it has been shown that LLMs are highly sensitive to irrelevant content Shi et al.
+For example, the general belief in ranking may not align with LLM’s preferences due to the self-attention mechanism of Transformers, which can focus on any token regardless of its position.
+Another aspect is selection; while humans can easily disregard irrelevant information in a context, it has been shown that LLMs are highly sensitive to irrelevant content Shi et al.
 (2023a).
-There are likely more gaps that further diverge from human behavior, such as repetition, which is generally considered detrimental for retrieval systems Xia et al.
-(2017), but maybe be useful to LLMs for relevance weighting.
+There likely exist more aspects that further diverge the LLM’s preference from that of humans, e.g., repetition.
+Repeated information is generally considered detrimental for retrieval systems Xia et al.
+(2017), but it may be useful to LLMs for weighting the relevance of context items.
 We empirically investigate this preference gap, focusing specifically on ranking and selection.
 As shown in Fig.
-1, the performance varied by around 1% when we randomized the top-5 retrieved passages [^1] We note that this is different from Liu et al.
-(2023), where a “loss in the middle” phenomenon is observed, meaning that the model are better at using the relevant information at the beginning or end of its input context.
-This is probably due to different number of passages used.
-In Liu et al.
-(2023), 20 total documents are employed, whereas we used 5 passages.
-Regardless, this does not affect the conclusion that there exists a preference gap between the retrievers and LLMs which needs to be bridged..
-However, the variation exceeded 5% when selecting the top-1 passage.
-This indicates the general belief in the importance of ranking in retrieval does not apply to LLMs , indicating that selection could be more crucial.
-This finding confirms the existence of the preference gap between the retrievers and LLMs, and highlights the importance of bridging this preference gap to enhance RAG.
-This is a crucial insight that has not been revealed before, which can guide RAG designers towards what should be achieved in the model.
-To design a RAG model, existing work has tried to finetune the LLMs to align with the retriever’s preference, or to finetune the retriever to align with the LLM’s preference.
-However, since they primarily focuses on ranking, their approaches are sub-optimal for bridging the preference gap.
-Fine-tuning retrievers can only achieve re-ranking, and fine-tuning LLMs is often impractical, as many are only accessible through APIs.
-To better bridge the gap beyond ranking, we propose a framework called BGM (Bridging the Gap between retrievers and LLMs).
-We propose to keep the retriever and LLM fixed, and train a bridge model in between.
-The bridge model aims to transform the retrieved information into a format that LLMs can work with effectively.
-Since the preference gap is present in both ranking and selection, and potentially extends to aspects such as repetition, in contrast to the conventional scoring format, we structure the bridge model as a sequence-to-sequence (seq2seq) model.
-This approach allows it to not only rerank but also dynamically select passages for each query, and potentially employ more advanced strategies like repetition.
-In a typical RAG, there is no ground truth relevance label for what should be retrieved but only ground truth label for the downstream task.
-Existing supervised learning (SL) approaches use the supervision provided by the LLM, such as the perplexity of downstream tasks.
-However, this is ineffective due to sparse supervision as it is nearly impossible to feed all possible retrieved sequences into the LLM to obtain supervision.
-Even worse, such SL relies on the intermediate relevance label and is not end-to-end training on the downstream tasks.
-To address these issues, we employed reinforcement learning (RL) on the SL trained bridge model, where the downstream performance metrics are used as the reward and the bridge model are regarded as a policy model.
-This framework, chaining SL and RL, provides increased supervision from the downstream task.
-It also offers the model the flexibility to explore more advanced strategies, such as repetition, in forming the optimal passage sequence.
+1, when we randomize the ordering of top-5 retrieved items (in our case, passages), the performance of RAG only varies by around 1%  [^1] We note that this finding is different from Liu et al.
+(2023), where a “loss in the middle” phenomenon is observed, meaning that the LLM is better at using the relevant information at the beginning or end of its input context.
+This is probably due to different numbers of retrieved items.
+Liu et al.
+(2023) used 20 documents, whereas we use 5 passages.
+Regardless, both findings indicate that there exists a preference gap between the retrievers and LLMs, which we aim to bridge..
+However, the variation exceeds 5% when the LLM is only presented with the top-1 passage under each order (therefore it encounters different selections of information).
+This indicates the general belief in ranking does not always apply to LLMs, and that the selection of information could be more crucial.
+This finding confirms the existence of the preference gap between retrievers and LLMs, and it is critical to bridge this preference gap to enhance the performance of RAG.
+To the best of our knowledge, this is a novel insight that may guide future designs of RAG systems.
+Existing work has tried to finetune the LLMs to align with the retriever or adjust the retriever to align with the LLM.
+However, finetuning LLMs, especially at the scale of GPT-4 or Palm 2, is often expensive.
+Similarly, it is difficult to adjust production-level retrievers such as Google or Bing.
+Even when the retriever is adjustable, existing efforts often focus on re-ranking the retrieved results and fail to address other aspects of preference such as selection or repetition.
+Instead, we propose a novel and practical framework called BGM (Bridging the Gap between retrievers and LLMs), which keeps the retriever and LLM fixed and trains a bridge model in between.
+The bridge model aims to transform the retrieved information into a format that LLMs prefers and can effectively work with.
+Without loss of generality, we structure the bridge model as a sequence-to-sequence (seq2seq) model, which allows dynamically selecting items, re-ranking them, and potentially broader operations like repeating some of them in the retrieval-augmented prompt.
+Training such a bridge model is challenging as there are usually no ground truth labels on ideal item sequences for retrieval augmentation.
+Existing work has tried to derive supervisory signals for ranking from RAG’s downstream task performance, such as perplexity distillation  Izacard et al.
+(2022).
+Nevertheless, these methods only provide pointwise supervisory signals for each item independently.
+Directly applying the same idea to obtain sequential supervision is infeasible, since it would require feeding all possible item sequences into the LLM to obtain perplexity.
+We developed a greedy search approach to solve this problem (Sec.
+4.1).
+Moreover, we find sequential supervision can be too sparse to effectively train such a seq2seq model (Table 5).
+To address this issue, we employed reinforcement learning (RL) on the SL trained bridge model, regarding the downstream task performance as the reward and the bridge model as a policy model.
+Chaining SL and RL provides increased supervision from the downstream task.
+It also offers the flexibility to explore more advanced strategies, such as repetition, in forming the optimal passage sequence.
 Our experiments reveal that BGM can enhance the performance of various downstream tasks, such as Question Answering (QA) and personalized generation, across a spectrum of datasets, from public QA and amazon reviews to private email conversations.
-Notably, the modified passages retrieved by BGM surpass the performance of strong retrievers and reranking baseline models.
-This underscores the significance and promise of the bridging approach in the realm of RAG.
+Notably, the modified passages retrieved by BGM surpass the performance of strong retrievers and baseline reranking models.
+This underscores the significance and promise of the “bridge” approach in the realm of RAG.
 In summary, our contributions can be summarized as follows:
-  • We empirically establish the existence of the preference gap between retrievers and LLMs, and introduce BGM, which designed to address this preference gap.
+  • We empirically establish the existence of the preference gap between retrievers and LLMs, and introduce BGM to address this gap.
   • We propose a seq2seq bridge model to jointly accomplish reranking and selection, adapting the retrieved information to be LLM-friendly.
 We employ a SL and RL training scheme to optimize this adaptation process.
-  • The evaluation of BGM comprises diverse datasets, including QA and text generation, including both publicly available and personal information sources.
-The comprehensive evaluation underscores the effectiveness of BGM in bridging the preference gap and improving performance in downstream tasks.
-Figure 2: Differing from previous RAG that update either LLMs, retrievers, or both, BGM not only treats the LLM as a black box but also freezes the retrieval.
-Only a lightweight bridge model (a Seq2seq LM) is tunable in order to adapt the retrieved information to satisfy the LLM’s preference.
-This makes BGM applicable to large LLMs, which are often served via APIs and also avoids expensive updates to the retriever.
+  • We evaluate BGM with diverse tasks, including QA and text generation, with publicly available and personalized datasets.
+The evaluation underscores the effectiveness of BGM in bridging the preference gap and improving RAG performance in downstream tasks.
+Figure 2: Differing from previous RAGs that update LLMs, retrievers, or both, BGM connects a frozen LLM and a frozen retriever through a lightweight bridge model which adapts the retrieved information to the LLM’s preference.
+This makes BGM applicable to “large” LMs and off-the-shelf retrievers.
 2 Related Work
 Retrieval-augmented Generation (RAG).
-Augmenting LLMs with relevant information retrieved from various knowledge sources has shown effective in improving performance across numerous NLP tasks, including language modeling Borgeaud et al.
+Augmenting LLMs with relevant information retrieved from various knowledge sources is proven effective across numerous NLP tasks, including language modeling Borgeaud et al.
 (2022); Khandelwal et al.
@@
 (2020).
-Specifically, RAG utilizes input as a query and comprises two main components: (1) retriever retrieves a set of items from a large corpus.
-Note that the retrieval units vary across different works, including documents, passages, or even tokens.
-In this study, we focus on retrieving passages; and (2) LLM incorporate the retrieved passages as additional information, integrating them into the input context for making final predictions.
-A fundamental question in this process arises regarding the differing preferences between LLMs and existing retrievers, with LLMs performing optimally only when their preferences are satisfied.
-Bridging the gap between these differing preferences is crucial.
-Depending on which components are subject to updates, this challenge can be categorized into three families
+Specifically, RAG utilizes input as a query and comprises two main components: (1) a retriever retrieves a set of items from a side corpus.
+Particular items may vary across different tasks, including documents, passages, or even tokens.
+In this study, we focus on retrieving passages; and (2) a LLM incorporates the retrieved items, as additional information in the input context, and makes final predictions.
+A fundamental question in this process arises regarding the disparate preferences between LLMs and retrievers, as LLMs performing optimally only when their preferences are satisfied.
+Bridging the preference gap is crucial.
+Depending on which components are subject to updates, this challenge can be categorized into three families.
 Finetuning retrievers and LLMs jointly.
-This is the most widely used and conventional setting of RAG Izacard et al.
+This is the most widely used setting of RAG Izacard et al.
 (2022); Khandelwal et al.
@@
 (2020).
-However, they are mostly less than 1B so may nor regard as “large” LM.
+However, most prior work is based on relative small LMs ($<$ 1B).
 For example, Altas Izacard et al.
@@
 (2020) uses a tunnable query encoder and DPR Karpukhin et al.
-(2020) as retriever, BART as LLM, and design an end-to-end training and inference schema to train the query encoder and the LLM.
+(2020) as retriever, BART as LLM, and design an end-to-end schema to train the query encoder and the LLM.
 Finetuning LLMs only.
@@
 Although above systems have shown improves results, they are not always applicable in practice.
-Many LLMs such as GPT-3 and Codex are not open-sourced due to commercial considerations and are only available as black-box APIs.
-Through these APIs, users can send queries and receive responses.
+Many LLMs (especially larger ones) are only available as black-box APIs.
 Given this constraint, a natural approach is to only update the retrievers to ensure they can retrieve passages that are more compatible with LLMs.
@@
 However, this family of models is incapable of performing any sample-level selection and can only choose top passages by setting a fixed threshold.
-Unlike the existing three families, BGM works on LLM (> 11B) and directly working on bridging the preference gap, without finetuning LLMs or retrievers but a bridge model in between (Fig. 2).
+Unlike the existing approaches, BGM does not fine-tune the LLM or the retriever and instead employs a bridge model in between (Fig. 2).
 RL for information retrieval.
@@
 Typically, an IR task would be structured to determine which document to select for a ranking position, using ranking metrics such as DCG as the reward.
-This approach had advantages, including using the ranking position directly as supervision and optimizing the IR measure directly.
-However, none of these existing works explored the application of RL in the context of RAG.
+None of these existing studies explore the application of RL in the context of RAG.
 In the LLM era, RL has been used in query rewriting for retrieval Wu et al.
@@
 Bacciu et al.
-(2023) suggest using RL to fine-tune retriever in the RAG context in their opinion paper, but it does not conduct any experiments.
-More crucially, their work fails to recognize the importance of bridging the gap between retrievers and LLMs, nor does it specify what the bridge model should be.
+(2023) suggest using RL to fine-tune retriever in the RAG context in their opinion paper, not supported by experiments.
+Their work does not recognize the importance of bridging the gap between retrievers and LLMs, nor does it specify what the bridge model should be.
 3 Problem Formulation
@@
 Given an input $x$, the retriever aims to retrieve a ranked list of passages from a corpus $D=\{d_{i}\}_{i=1}^{m}$ that are relevant to $x$.
-In this work, we assume a frozen dense retriever.
+In this work, we assume a typical scenario of employing a frozen dense retriever.
 Typically, a dual encoder architecture is applied, where an encoder is used to encode both the input context $x$ and the passage $d$.
@@
 $s(d,x)=\text{cos}(\bm{E}(d),\bm{E}(x)).$ (1)
-The top-k passages that have the highest similarity scores when compared with the input $x$ are retrieved in this step.
+The top-k passages that have the highest similarity scores when compared with the input $x$ are retrieved in this step,
 $(d^{\text{retr.}}_{j})_{j=1}^{k}=\text{Top-K}(\{s(d,x)\}_{i=1}^{m}).$ (2)
 Bridge Model for RAG.
-The retrieved top-K passages provide richer information about the original input $x$ and can potentially help the LLM to make a better prediction on downstream tasks.
-However, since there is preference gap, we propose a bridge model $\bm{B}$ to adapt the retrieved passages to LLM-friendly passages.
+The retrieved top-K passages provide richer information about the original input/query $x$ to help the LLM to make a better prediction on downstream tasks.
+A bridge model $\bm{B}$ adapts the retrieved passages to a sequence of passages that is LLM-friendly.
 As mentioned in the Sec.
 1, the bridge model is a seq2seq model.
-It takes as input all the retrieved passages $(d^{\text{retr.}}_{j})_{j=1}^{k}$ and output a subset of adapted passages $(d^{\text{bdr.}}_{j})_{j=1}^{n}$.
-This format is advantageous because the seq2seq model automatically considers ranking by generating the next token based on the preceding one and selection by placing the end-of-sentence token in the appropriate position.
-Note that $n$ may or may not equal to $k$ due to selection.
-$(d^{\text{bdr.}}_{j})_{j=1}^{n}=\bm{B}((d^{\text{retr.}}_{j})_{j=1}^{k}).$ (3)
-It is important to note that in practice, the bridge model’s input includes the query, passage IDs, and passage content, while the output consists of passage IDs.
-We then convert the obtained passage IDs to their corresponding documents when we use them in RL (Sec.
-4.2).
-Additionally, the passages in passage sequence (e.g., $(d^{\text{retr.}}_{j})_{j=1}^{n}$) are concatenated into a long sequence before being fed to the network.
-Retrieval-augmented Generation with bridge model.
-Given the adapted passages from the bridge model as context, $(d^{\text{bdr.}}_{j})_{j=1}^{n}$, we concatenate them with the input $x$, and fed the resulting long sequence into the LLM to obtain the output for downstream tasks.
+It takes all the retrieved passages $(d^{\text{retr.}}_{j})_{j=1}^{k}$ as well as the query $x$ as input, and outputs the adapted passages $(d^{\text{bdr.}}_{j})_{j=1}^{n}$,
+$(d^{\text{bdr.}}_{j})_{j=1}^{n}=\bm{B}\left(x,(d^{\text{retr.}}_{j})_{j=1}^{k}\right).$ (3)
+This formulation is general enough as the seq2seq model automatically considers ranking by generating the next token based on the preceding one, selection by placing the end-of-sentence token in the appropriate position, and repetion by generating the same passage ID (as explained in the following paragraph).
+Note that $n$ may be smaller or larger than $k$ due to selection and repetition.
+Before concatenating the query and passages as bridge model’s input, we prepend each passage with a unique sentinel token as its passage ID, e.g., $[\text{query}][\text{id}_{1}]d^{\text{retr.}}_{1}[\text{id}_{2}]d^{\text{retr.}}_{2}$.
+In this way, the model only needs to generate the passage IDs instead of the actual passage content, which is much more efficient and avoids making unfaithful changes to the retrieved passages.
+We then convert the obtained passage IDs to the corresponding passages for downstream processing.
+Retrieval-augmented generation with bridge.
+We concatenate adapted passages from the bridge model, $(d^{\text{bdr.}}_{j})_{j=1}^{n}$, with the input $x$, and fed the resulting long sequence into the LLM as context to obtain the output for downstream tasks.
 Figure 3: Illustrating the training of BGM: Step 1: First, we prepare the silver passage sequence for supervised learning (SL) through a greedy search on the retrieved passages.
-Step 2: This silver passage sequence is then used for the supervised training of the bridge model.
-The bridge model is a seq2seq model that takes the query, passage ID, and passage content as input and outputs the passage IDs.
-Step 3: Finally, the trained bridge model is treated as a policy model and is further trained using reinforcement learning
+Step 2: These silver passage sequences are then used for the supervised training of the bridge model.
+The bridge model is a seq2seq model that takes the query and passages with prepended passage IDs as input and outputs the passage IDs.
+Step 3: Finally, the SL-trained bridge model is treated as a policy model and is further trained using reinforcement learning.
 4 Training the Bridge Model
@@
 3, we format the bridge model as seq2seq model.
-However, it is challenging to effectively train it as there is no ground truth retrieved passages and LLM is assumed to be API without gradient.
-To address the issue, we propose to chain supervised learning (SL) and reinforcement learning (RL), where SL aims to reduce the search space of RL and provides a reasonably good initial model that does ranking and selection.
-Meanwhile, RL aims to optimize the policy model, i.e., bridge model, against the reward.
-4.1 Supervised Learning
-To conduct SL, the ground-truth sequence is required for each query.
-In a typical RAG, the ground truth relevance passage for the query is not provided, let alone the ground-truth sequence.
-To address this, we propose to synthesis silver passage sequence by selecting only the useful passage.
-This is done by greedy search that incrementally selects the next passage that can maximize the downstream task performance.
-Synthesising silver passage sequence using greedy search.
-Assume that the downstream task performance for a given passage sequence of any length, used as context, is denoted as $R(\cdot)$.
-The empty passage sequence, i.e., no retrieved passage, denoted as $d^{\text{retr.}}_{\varnothing}$, has downstream task performance $R(d^{\text{retr.}}_{\varnothing})$.
-Starting from $d^{\text{retr.}}_{\varnothing}$ and empty silver passage sequence $d^{\text{silv.}}$, we try to add (i.e., concatenate) the candidate passage that can improve most to the silver passage sequence, until no improvement can be made.
-Algorithm 1 shows the pseudo-code for synthesising silver passage sequence.
-Input: $(d^{\text{bdr.}}_{j})_{j=1}^{n}$, $R(\cdot)$
+As mentioned in Sec.
+1, its training is challenging due to the lack of supervision for passage sequences, the infeasibility of directly applying existing methods that provide only point-wise supervision, and the sparsity of sequential supervision.
+For effective training, we propose to chain supervised learning (SL) and reinforcement learning (RL), where SL aims to reduce the search space of RL and provides a reasonably good initial model that does ranking and selection.
+RL aims to optimize the policy model, i.e., bridge model, for the downstream task.
+Fig.
+3 shows an overview.
+Input: $(d^{\text{retr.}}_{j})_{j=1}^{n}$, $R(\cdot)$
 Output: $(d_{j}^{\text{silv.}})_{j=1}^{s}$
-1 $d^{\text{silv.}}=()$;
-2 $silv=R(d^{\text{retr.}}_{\varnothing})$;
-3 $prev=0$;
-4 $best=\varnothing$;
+1 $d^{\text{silv.}}\leftarrow()$;
+2 $R^{\text{silv.}}\leftarrow R(d^{\text{silv.}})$;
+3 $R^{\text{best}}\leftarrow-\infty$;
 4 while true do
-6       for $d\leftarrow d^{\text{retr.}}_{1}$ to $d^{\text{retr.}}_{n}$ where $d$ not in $d^{\text{silv.}}$ do
-7             $seq$ = concat($d$, $d^{\text{silv.}}$);
-8             $cur$ = $R(seq)$;
-9             if (cur > silv) and (cur > prev) then
-10                   $best$ = $d$;
-11                   $prev$ = $cur$;
+5       for $c\leftarrow d^{\text{retr.}}_{1}$ to $d^{\text{retr.}}_{n}$ where $c\not\in d^{\text{silv.}}$ do
+6             $d^{\prime}\leftarrow\mathtt{add\_to\_seq}(d^{\text{silv.}},c)$;
+7             $R^{\text{cur}}\leftarrow R(d^{\prime})$;
+8             if $R^{\text{cur}}>R^{\text{best}}$ then
+9                   $c^{\text{best}}\leftarrow c$;
+10                   $R^{\text{best}}\leftarrow R^{\text{cur}}$;
+11
 12
-13
-14      if best != $\varnothing$ then
-15             add $best$ to the end of $d^{\text{silv.}}$;
-16             $silv=prev$;
-17             $prev=0$;
-18             $best=\varnothing$
+13      if $R^{\text{best}}>R^{\text{silv.}}$ then
+14             $d^{\text{silv.}}\leftarrow\mathtt{add\_to\_seq}(d^{\text{silv.}},c^{\text{best}})$;
+15             $R^{\text{silv}}\leftarrow R^{\text{best}}$;
+16             $R^{\text{best}}\leftarrow-\infty$;
+17
 18      else
 19             break;
+20
 21
-22
-Algorithm 1 Synthesis Silver Passage Sequence
-Training. Training with the silver passage sequence is straightforward, achieved by applying cross-entropy loss.
+Algorithm 1 Synthesis SPS
+4.1 Supervised Learning
+To conduct SL, the ground-truth passage sequence is required for each query.
+Existing approaches, which focus on obtaining the relevance score of each passage (see Section 1), are not applicable as we need to determine which combination of passages is most effective for the downstream tasks.
+To address this, we propose to synthesis silver passage sequence (SPS) by selecting only the useful passages.
+This is done by greedy search that incrementally selects the next passage that can maximize the downstream task performance.
+Synthesising SPS using greedy search.
+We denote the downstream task performance when using a given passage sequence for RAG as $R(\cdot)$.
+For the edge case where the passage sequence is empty, $R(\varnothing)$ simply denotes the task performance without retrieval augmentation (i.e., no passage used).
+We start from an empty SPS $d^{\text{silv.}}=\varnothing$, and iteratively add the next best candidate passage to the sequence, measured based on the resulted task performance $R(d^{\text{silv.}})$.
+We stop until no improvement can be made to $R(d^{\text{silv.}})$.
+Algorithm 1 shows the pseudo-code for synthesising SPS.
+The training with the SPS is achieved by applying cross-entropy loss.
 4.2 Reinforcement Learning
-Although supervised learning can already help training the bridge model, it is still ineffective (we can see using SL alone results in mixed performance compared to using a retriever alone in Table 5).
-This is attributed to sparse supervision and the lack of end-to-end training on downstream results.
-To address these issues, we apply RL to further train the bridge model.
-RL does not limit the possible passage sequences (we only consider permutations or deletions in the silver passage sequence in SL.
-However, the optimal passage sequence might require more complex manipulations, such as repetition, which RL can accommodate) and give more supervision to find the best passages sequence.
-Using the performance of downstream task as the reward, the bridge model is trained in an end-to-end manner on the downstream task.
-Specifically, our task can be formulated as a RL problem as follows:
-Reward is the performance of tasks against their ground-truth labels as we focus on labeled downstream tasks.
+Although SL can already help training the bridge model, it is still ineffective – we observe using SL alone results in mixed performance (see Table 5).
+This is attributed to sparse supervision (see Sec.
+1) and the lack of end-to-end training on downstream results.
+To address these issues, we apply RL to continue the training of the bridge model.
+In SL, we only consider permutations or deletions in the SPS, while RL can accommodate more complex manipulations that an optimal passage sequence might require, such as repetition.
+Additionally, RL provides enhanced supervision beyond the silver sequences through the reward from sampled passage sequences.
+Using the performance of downstream task as the reward, the bridge model is trained in an end-to-end manner.
+Specifically, our task can be formulated as an RL problem – Reward is the performance of the downstream task, usually measured based on certain ground-truth labels, e.g., Extract Match or BLEU.
 Policy model is the bridge model that need to be trained.
-Action Space is limitted to Passage IDs as we are interested in adapting the retrieved passages
-Training. The reward objective can be optimized by any off-the-shelf RL algorithm, e.g., proximal policy optimization (PPO).
+Action space is restricted to passage IDs (Sec.
+3) as we are interested in organizing rather than revising the retrieved passages.
+In training, the reward objective can be optimized by any off-the-shelf RL algorithm, e.g., proximal policy optimization (PPO).
 5 Experiments
@@
 Datasets.
-We consider four datasets, ranging from popular QA datasets to less investigated personalized ones.
+We consider four datasets, ranging from popular QA datasets to personalized generation datasets.
 We also include one dataset that contains private email conversations (Avocado Email), which is unlikely to be included in the LLM’s pre-training datasets.
@@
 Open-domain QA.
-We conduct evaluations on two open-domain QA: Natural Questions (NQ) Kwiatkowski et al.
+We conduct evaluations on two open-domain QA datasets: Natural Questions (NQ) Kwiatkowski et al.
 (2019) and HotpotQA Yang et al.
@@
 HotpotQA is a multi-hop QA dataset which requires findings and reasoning over multiple passages to answer the question.
-The candidate passages are retrieved from WikiPedia pages
+The candidate passages are retrieved from Wikipedia pages.
 Personalized Generation.
@@
 (2023) to construct the personalized generation datasets.
-Specifically, suppose a user is writing a document, which we call the current document.
-Given the immediate context $x$ and the user’s personal context, we aim the finish the document as if the user had completed it.
-The immediate context $x$ or the input to the personalized writing task, is defined as the title and the start of the current document.
-The candidate passages are retrieved from documents authored by this user in the past.
+In this task, the query includes the start and necessary properties (e.g., title) of a document authored by a user.
+The objective is to complete the document as if it were finished by the same user.
+The candidate passages for this task are retrieved from documents previously authored by this user.
 This includes datasets from two domains: Avocado Email (Email) Oard et al.
@@
 (2019).
+An example from Book is given in Fig.
+4.
 Figure 4: An example from the Book dataset.
-The query consists of the instruction, the document (i.e., review for Book) title, document product (there will be no product information for Email), document start.
-The target is to complete the text given the query.
+The query consists of the instruction, title, product (there will be no product information for Email) and document start.
+The goal of the task is to generate the remaining part of the document.
 Baselines.
@@
 This is a variant from Izacard et al.
-(2022) where we substitute the decoder with our LLM and apply point-wise scoring reranking.
-This approach can be regarded as utilizing a reranker as a bridge model, but lacking the capability for dynamic selection; and (4) Additionally, we include a non-retrieval baseline, Naive, where no retrieval is conducted to aid in the generation process.
-#Training #Val. #Test Avg. # words
+(2022) where we substitute the decoder with our LLM and apply perplexity distillation.
+This approach can be regarded as utilizing a reranker as a bridge model, but lacking the capability for dynamic selection; and (4) Additionally, we include a non-retrieval baseline, Naive, where no retrieval augmentation in the generation process.
+#Training #Val. #Test Avg. #Tokens
 NQ 79,168 8,757 3,610 517.82
@@
 Table 1: Statistic of the 4 datasets.
+“Avg.
+# tokens” indicates the average number of tokens in the prompt (which includes the query and retrieved passages).
+We ensure that the length does not exceed the maximum length allowed by the LLM.
 5.2 LLM and Hyperparamters
@@
 (2020b) model as our bridge model for most experiments.
-In the supervised learning stage, the T5-11B model is fine-tuned with a base learning rate of 0.001.
+In the supervised learning stage, the T5-XXL model is fine-tuned with a base learning rate of 0.001.
 A linear warmup scheduler is used for the first 1,000 training steps.
@@
 Decoding is performed using beam search with a beam size of 4.
-We use the PaLM2 model Anil et al.
+We use the PaLM2-S model Anil et al.
 (2023), a new state-of-the-art LLM, as our LLM.
 It adopts temperature sampling as the decoding strategy.
-The parameters of PaLM2 are frozen, and we set the temperature to 0 to make the output deterministic.
-We use Exact-Match (EM) and BLEU as the metric to select the best prompt among prompt variants.
+The parameters of PaLM2-S are frozen, and we set the temperature to 0 to make the output deterministic.
+We use Exact-Match (EM) and BLEU metrics, depending on the dataset.
 It is also used as the reward to train the bridge model in the RL stage.
-The “Top-K” in Eq.
+The $K$ in “Top-K” in Eq.
 2 is set to 5.
@@
 Metric EM EM BLEU BLEU
-Naïve 33.07 28.01 5.57 11.5
+Naive 33.07 28.01 5.57 11.5
 Random 43.71 26.10 8.55 8.61
@@
 BGM 45.37 35.64 10.42 12.07
-Table 2: Performance of all 4 datasets
-Superior of BGM. Table 2 reports the overall performance. We can see that
-(1) The proposed BGM outperforms all the 4 baselines in all 4 datasets. This clearly indicates BGM is effective in adapting the retrieved passages.
+Table 2: Performance of all 4 datasets.
+Superiority of BGM. Table 2 reports the overall performance. We can see that
+(1) The proposed BGM outperforms all the 4 baselines in all 4 datasets.
+This clearly indicates BGM is effective in adapting the retrieved passages.
+It’s noteworthy that Random and GTR perform similarly, suggesting that ranking has less impact than selection of passages does.
 (2) Compared to the Naive approach, BGM shows significant improvement overall.
-However, an exception is observed in the Book dataset, where BGM’s improvement is more modest.
-This suggests that the LLM already possesses a substantial amount of relevant knowledge, likely due to the inclusion of Amazon reviews (which the Book dataset comprises) in the LLM’s pre-training data.
-Consequently, retrieval is not always essential in this context.
+An exception is observed in the Book dataset, where the improvement of BGM is less pronounced.
+This suggests that retrieval is not always essential in this context.
 This also explains why the Naive approach outperforms other baselines (except BGM) on the Book dataset.
-It’s important to note that BGM still manages to show improvement, albeit to a lesser extent, in this scenario
+It’s important to note that BGM still manages to show improvement by dynamically deciding whether and how many passages to select.
+For example, in the Book dataset, for around half of the samples (approximately 25,000), BGM selects no passages for augmentation.
 (3) Compared to the GTR approach, BGM demonstrates substantial improvement.
-The NQ dataset shows the least improvement, as most instances only require one retrieved passage, which both GTR and BGM can successfully include.
-On the other hand, the HotpotQA dataset shows the most significant improvement, indicating that HotpotQA may be more sensitive to irrelevant passages.
-It’s noteworthy that Random and GTR perform similarly, while GTR shows improvement over Naive (with the exception of the Book dataset).
-This suggests that the ranking itself does not significantly impact performance, but the selection of passages does
-(4) Compared to the PSR approach, BGM once again demonstrates significant improvement.
+Compared to HotpotQA, NQ shows smaller improvement, as most instances need only one passage, which both GTR and BGM can successfully include.
+Conversely, the HotpotQA shows a more substantial improvement, indicating that HotpotQA may be more sensitive to irrelevant passages.
+(4) Compared to the PSR approach, BGM again demonstrates significant improvement.
 This indicates that pure reranking alone is not sufficient for the bridge model.
 Selection must also be taken into account.
-It’s notable that PSR performs similarly to GTR, further suggesting that reranking alone does not greatly affect performance
+It’s notable that PSR performs similarly to GTR, further suggesting that reranking alone has limited impact.
 Model NQ HotpotQA Email Book
@@
 Naive 33.07 28.01 5.57 11.5
-PSR 43.6 25.51 9.08 9.14
+PSR 43.60 25.51 9.08 9.14
 PSR (Top1) 42.02 32.69 7.28 11.53
@@
 BGM 45.37 35.64 10.42 12.07
-Table 3: Ablation results - threshold the PSR
+Table 3: Ablation - threshold the PSR.
 Silver Data NQ HotpotQA Email Book
 Metric EM EM BLEU BLEU
-GTR 43.79 25.8 9.76 8.75
+GTR 43.79 25.80 9.76 8.75
 PSR 43.68 29.73 10.1 10.35
 Greedy (BGM) 45.37 35.64 10.42 12.07
-Table 4: Ablation results - different silver passage sequence for SL
+Table 4: Ablation - different SPS for SL.
 Model NQ HotpotQA Email Book
@@
 BGM 45.37 35.64 10.42 12.07
-Table 5: Ablation results - SL only
+Table 5: Ablation - SL only.
 Model NQ HotpotQA Email Book
@@
 FLAN-T5-XL 44.87 35.41 9.64 10.7
-FLAN-T5-XXL (BGM) 45.37 35.64 10.42 12.07
-Table 6: Ablation results - different bridge model size
+FLAN-T5-XXL 45.37 35.64 10.42 12.07
+Table 6: Ablation - different BGM bridge model size.
 Model Palm2-XXS Palm2-S
-Data NQ HotpotQA Email Book
-Metric EM EM BLEU BLEU
+Data NQ HotpotQA NQ HotpotQA
+Metric EM EM EM EM
 Naive 12.13 14.57 33.07 28.01
-Random 31.19 24.41 43.71 26.1
-GTR 31.91 23.17 43.79 25.8
+Random 31.19 24.41 43.71 26.10
+GTR 31.91 23.17 43.79 25.80
 PSR 32.04 22.53 43.60 25.51
 BGM 39.88 28.69 45.37 35.64
-Table 7: Ablation results - different LLM size
+Table 7: Ablation - different LLM size.
 Model NQ HotpotQA Email Book
@@
 Test on Palm2-S
-BGM 45.37 35.64 10.42 12.07
+BGM (in-domain) 45.37 35.64 10.42 12.07
 BGM
-(Train on NQ)
+(Trained on NQ)
 33.42 5.66 11.22
 BGM
-(Train on Email)
+(Trained on Email)
 35.59 27.98 11.38
@@
 BGM
-(Train on Palm2-XXS)
+(Trained with Palm2-XXS)
 39.88 28.69
 BGM
-(Train on Palm2-S)
+(Trained with Palm2-S)
 30.63 24.55
-Table 8: Ablation results - the generability of BGM across various datasets and LLM sizes.
-Understanding BGM. We are interested in what happen when we optimize the BGM, whether BGM’s improvement is trival, and how some important setting affects the BGM.
-(1) Whether the selection can be achieved naively by manually threshold the PSR?
+Table 8: Ablation - the generability of BGM across various datasets and LLM sizes.
+Understanding BGM. Next, we study a few research questions about BGM and RAG.
+(1) Can we perform effectively passage selection for RAG by simply thresholding PSR?
 We conducted an ablation experiment using various thresholds for PSR, as shown in Table 3.
-The term “Top-K” in rows 3 to 6 refers to the selection of only the Top-K passages from PSR reranked passages.
-The results are mixed, but consistently inferior to those of BGM.
-This suggests that a naive manual threshold applied to the reranking model is insufficient to meet our objectives.
-Therefore, it is necessary to consider both reranking and dynamic selection simultaneously
-(2) How different silver passage sequence will affect the final performance?
-In Table 2, we employed greedy search passage sequences as silver passage sequence for SL (Sec.
+“Top-K” in rows 3 to 6 denotes selecting only the top-k passages from PSR reranked passages.
+The results vary but are consistently lower than BGM’s.
+This suggests that a naive manual threshold applied to the reranking model is insufficient to achieve the desired objectives.
+Therefore, it is necessary to consider both reranking and dynamic selection simultaneously.
+(2) How does different SPS (sliver passage sequence) used for SL affect BGM performance?
+In Table 2, we employed greedy search passage sequences as SPS for SL (Sec.
 4.1).
-However, lacking ground truth passage sequences, we also explored whether this approach is superior to others.
-Our ablation experiment in Table 4 involved using various types of silver passage sequence.
-The first row, labeled ’GTR’, indicates the use of GTR’s retrieved passages as silver passage sequence.
-Similarly, ’PSR’ refers to using PSR reranked passages as silver passage sequence.
-In the final row, we used greedy search passage sequences, representing the current version of BGM.
-The results demonstrate that the quality of silver passage sequence significantly affects downstream task performance.
-While using PSR as silver passage sequence shows improvement over GTR on three datasets, Greedy (BGM) further enhances performance and achieves the best results.
-Identifying potentially better silver passage sequence is left for future work.
-(3) How helpful RL is to the final performance?
+Here, we explore whether this approach is superior to others.
+Our ablation experiment in Table 4 involved using various types of SPS.
+The first row, labeled “GTR”, indicates the use of GTR’s retrieved passages as SPS.
+Similarly, “PSR” refers to using PSR reranked passages as SPS.
+In the final row, we used greedy search passage sequences, representing the proposed version of BGM.
+The results demonstrate that the quality of SPS significantly affects downstream task performance.
+While using PSR as SPS shows improvement over GTR on three datasets, Greedy (BGM) further enhances performance and achieves the best results.
+Identifying potentially better SPS is left for future work.
+(3) How helpful is RL to BGM?
 BGM integrates SL and RL, and we aim to assess the effectiveness of each component.
@@
 This suggests that SL alone is inadequate, highlighting the necessity of incorporating RL.
-It’s important to note that removing SL is not feasible, as this would result in an excessively large search space for RL to effectively learn.
-(4) How different size of bridge model affect the final performance?
-In Table 2, we utilized Flan-T5-XXL (11B) as the bridge model, which is already smaller than the LLM (24B).
-However, we are curious about the feasibility of using an even more lightweight language model (LM) as the bridge model.
+Note that removing SL is likely ineffective, leading to a poorly initialized policy model for RL with an excessively large search space.
+(4) How does the size of bridge model affect the final performance?
+In Table 2, we utilized Flan-T5-XXL (11B) as the bridge model, which is already smaller than the PaLM2-S.
+However, we are curious about the feasibility of using an even more lightweight LM as the bridge model.
 The ablation study shown in Table 6 presents the outcomes of employing bridge models of various sizes.
@@
 This demonstrates that a bridge model is beneficial even at a smaller scale, and that larger sizes lead to further improvements.
-(5) How different size of LLMs affect the final performance?
+(5) How does the size of LLMs affect the final performance?
 In Table 2, Palm2-S was used as the LLM.
 We are interested in evaluating the effectiveness of the bridge model with different sizes of LLMs.
-We conducted experiments using Palm2-XXS, as detailed in Table 7.
+We conducted experiments using Palm2-XXS in Table 7.
 It is important to note that the smaller LLM struggles with personalized generation datasets (i.e., Email and Book), resulting in BLEU scores lower than 1%.
@@
 This indicates that BGM is effective even with a smaller LLM.
-(6) Can a trained bridge model generalize to different datasets and LLMs?
-The effectiveness of the bridge model across various datasets and LLMs has been demonstrated.
-A more ambitious objective is to extend this performance to new contexts without additional training.
-We conducted experiments to investigate the generalizability of the bridge model.
-The upper section of Table 8 presents the data on generalizability across different datasets.
-Row 2 shows the performance of BGM (i.e., the bridge model) when trained exclusively on the NQ dataset and then tested on three other unseen datasets.
-Similarly, row 3 describes training BGM solely on the Email dataset and testing it on the other three.
-In all cases, the performance is inferior to that of BGM when both training and testing are conducted on the same datasets.
-This result is expected, as no specific techniques have been developed for dataset generalization, a topic we leave for future work.
+(6) Can a bridge model generalize to different datasets and LLMs?
+Our experiments above demonstrate the effectiveness of the bridge model when trained “in-domain” (i.e., trained and tested on the same dataset) using PaLM2-S as the LLM.
+A more ambitious goal is to extend this performance to new datasets or LLMs without extra training.
+We conducted ablation experiments to investigate bridge model generalizability: Table 8’s upper section shows the results across different datasets.
+Row 2 shows the performance of the bridge model when trained exclusively on the NQ dataset and then tested on three other unseen datasets.
+Similarly, row 3 shows training BGM solely on the Email dataset and testing it on the other three.
+In all cases, performance falls short of BGM’s when trained and tested on the same datasets.
+This is expected, given the lack of techniques for dataset generalization, a topic we leave for future work.
 In the lower section of Table 8, we present the results of BGM when tested on Palm2-XXS, but trained on Palm2-S.
-Comparing with results of BGM both trained and tested on Palm2-XXS (as shown in row 1 of the bottom section), the mismatch between LLM sizes leads to a significant decline in performance.
-This suggests that BGM’s ability to generalize across different LLM sizes is currently limited.
-Addressing this limitation is considered an important direction for future research.
+Comparing with results of BGM both trained and tested on Palm2-XXS (see row 1 of the bottom section), the mismatch between the training and testing LLMs leads to a significant decline in performance.
+This suggests that BGM’s ability to generalize across different LLMs is currently limited.
+Addressing this is considered an important direction for future research.
 (7) Case Studies.
-We provided examples of GTR, PSR, and BGM in Table 9 in the Appendix for the NQ dataset.
-For question I, both GTR and PSR yield the same incorrect output, even though their contexts have different orders.
+We provided examples of GTR, PSR, and BGM in Table 9 in Appendix for the NQ dataset.
+For question I, both GTR and PSR yield the same incorrect output, even though both include a relevant passage for RAG.
 Only BGM provides the correct answer, indicating that additional irrelevant context can be noisy and detrimental to RAG’s performance.
-In question II, unlike question I, GTR and PSR produce different answers.
-This suggests that the ranking can sometimes influence the results.
-While the specific mechanism by which ranking affects outcomes remains unclear (we regard it as a valuable topic for future research), it is evident that we need to consider ranking, as it may or may not affect performance.
-Again, BGM selects the most relevant passage and delivers the correct answer.
-In question III, none of the candidate passages contain the answer (they discuss FaZe Clan and the number of subscribers, but do not identify who has the most subscribers).
+In question II, none of the candidate passages contain the answer (they discuss FaZe Clan and the number of subscribers, but do not identify who has the most subscribers).
 GTR and PSR provide incorrect answers, as their additional context is unhelpful.
-In contrast, BGM opts not to select any passages and answers the question using its own parameters, resulting in the correct answer.
-This demonstrates that retrieval-augmented processes are not always necessary, and BGM is capable of handling such cases
+In contrast, BGM opts not to select any passages and answers the question using its own memory, resulting in the correct answer.
+This demonstrates that retrieval-augmented processes are not always necessary, and BGM is capable of handling such cases.
 6 Conclusion
-This paper highlights the importance of bridging the preference gap between retrievers and LLMs.
-It begins by demonstrating the existence of this gap, especially in terms of ranking and selection.
-This provides an important guidance about what should be achieved in a RAG system.
-It then suggests the development of a lightweight bridge model.
-This component is designed to adapt user-friendly passages retrieved by the retrievers into formats more suitable for LLMs.
-To accomplish these objectives, we introduce a novel system named BGM.
-BGM treats the bridging task as a seq2seq problem, allowing for simultaneous consideration of both ranking and selection.
-It is trained through a combination of supervised learning and reinforcement learning, facilitating end-to-end training of the bridge model based on the performance in downstream tasks and is capble even when the LLM is black-box (i.e., API).
-Extensive experiments have demonstrated BGM’s effectiveness.
+This paper demonstrates the need to bridge the preference gap between retrievers and LLMs, which has various levels of impact on ranking and selection in RAG systems.
+We propose a bridge model, BGM, to adapt the output of a frozen retriever for frozen LLMs, formatting the task as a seq2seq problem.
+BGM chains supervised and reinforcement learning, for dense supervisions and end-to-end training.
+Our extensive experiments validate BGM’s effectiveness.
+7 Limitations
+While effective, BGM has certain potential limitations.
+Firstly, BGM is limited in terms of generalization across datasets and domains, as shown in Sec.
+5.3.
+Secondly, the silver passage sequences for supervised learning is synthesized using greedy search.
+Although Sec.
+5.3 demonstrates its effectiveness, it is worth exploring other synthesizing approaches.
+We leave these explorations to the future work, as bridging the preference gap is an critical but new problem in RAG.
+8 Ethics Statement
+This paper proposes a novel bridge model that can effectively bridge the preference gap between retrievers and Large Language Models (LLMs).
+We do not anticipate any negative consequences for individuals as a result of this research.
+In the event of system failure, the primary outcome would be suboptimal performance of an LLM, and we do not foresee notable ethical implications.
+Appendix A Case Studies
+Table 9: Cases comparisons for GTR, PSR and BGM from NQ dataset.
+Question I
+Where does new crust come from in sea floor spreading
+Answer
+“basaltic magma” or “volcanic activity”
+GTR Context
+Given the passage titles and contexts below:
+Title: Seafloor spreading;
+Context: Seafloor spreading Seafloor spreading is a process that occurs at mid-ocean ridges, where new oceanic crust is formed through volcanic activity and then gradually moves away from the ridge.
+Earlier theories (e.g.
+by Alfred Wegener and Alexander du Toit) of continental drift postulated that continents "ploughed" through the sea.
+The idea that the seafloor itself moves (and also carries the continents with it) as it expands from a central axis was proposed by Harry Hess from Princeton University in the 1960s.
+The theory is well accepted now, and the phenomenon is known to be caused by convection currents in the;
+Title: Oceanic crust;
+Context: frozen in the basalt.
+A symmetrical pattern of positive and negative magnetic lines emanates from the mid-ocean ridge.
+New rock is formed by magma at the mid-ocean ridges, and the ocean floor spreads out from this point.
+When the magma cools to form rock, its magnetic polarity is aligned with the then-current positions of the magnetic poles of the Earth.
+New magma then forces the older cooled magma away from the ridge.
+This process results in parallel sections of oceanic crust of alternating magnetic polarity.
+Oceanic crust Oceanic crust is the uppermost layer of the oceanic portion of a tectonic;
+Title: Seafloor spreading;
+[…]
+Title: Seafloor spreading;
+Context: As new seafloor forms and spreads apart from the mid-ocean ridge it slowly cools over time.
+Older seafloor is therefore colder than new seafloor, and older oceanic basins deeper than new oceanic basins due to isostasy.
+If the diameter of the earth remains relatively constant despite the production of new crust, a mechanism must exist by which crust is also destroyed.
+The destruction of oceanic crust occurs at subduction zones where oceanic crust is forced under either continental crust or oceanic crust.
+Today, the Atlantic basin is actively spreading at the Mid-Atlantic Ridge.
+Only a small portion of the oceanic;
+Title: Seafloor spreading;
+Context: other areas), material from the upper mantle rises through the faults between oceanic plates to form new crust as the plates move away from each other, a phenomenon first observed as continental drift.
+When Alfred Wegener first presented a hypothesis of continental drift in 1912, he suggested that continents ploughed through the ocean crust.
+This was impossible: oceanic crust is both more dense and more rigid than continental crust.
+Accordingly, Wegener’s theory wasn’t taken very seriously, especially in the United States.
+Since then, it has been shown that the motion of the continents is linked to seafloor spreading by the;
+What is the answer of the following question:
+where does new crust come from in sea floor spreading
+GTR Prediction: the upper mantle
+PSR Context
+Given the passage titles and contexts below:
+Title: Seafloor spreading;
+Context: Seafloor spreading Seafloor spreading is a process that occurs at mid-ocean ridges, where new oceanic crust is formed through volcanic activity and then gradually moves away from the ridge.
+Earlier theories (e.g.
+by Alfred Wegener and Alexander du Toit) of continental drift postulated that continents "ploughed" through the sea.
+The idea that the seafloor itself moves (and also carries the continents with it) as it expands from a central axis was proposed by Harry Hess from Princeton University in the 1960s.
+The theory is well accepted now, and the phenomenon is known to be caused by convection currents in the; Title: Seafloor spreading;
+[…]
+Context: other areas), material from the upper mantle rises through the faults between oceanic plates to form new crust as the plates move away from each other, a phenomenon first observed as continental drift.
+When Alfred Wegener first presented a hypothesis of continental drift in 1912, he suggested that continents ploughed through the ocean crust.
+This was impossible: oceanic crust is both more dense and more rigid than continental crust.
+Accordingly, Wegener’s theory wasn’t taken very seriously, especially in the United States.
+Since then, it has been shown that the motion of the continents is linked to seafloor spreading by the;
+Title: Oceanic crust;
+Context: frozen in the basalt.
+A symmetrical pattern of positive and negative magnetic lines emanates from the mid-ocean ridge.
+New rock is formed by magma at the mid-ocean ridges, and the ocean floor spreads out from this point.
+When the magma cools to form rock, its magnetic polarity is aligned with the then-current positions of the magnetic poles of the Earth.
+New magma then forces the older cooled magma away from the ridge.
+This process results in parallel sections of oceanic crust of alternating magnetic polarity.
+Oceanic crust Oceanic crust is the uppermost layer of the oceanic portion of a tectonic;
+Title: Seafloor spreading;
+Context: As new seafloor forms and spreads apart from the mid-ocean ridge it slowly cools over time.
+Older seafloor is therefore colder than new seafloor, and older oceanic basins deeper than new oceanic basins due to isostasy.
+If the diameter of the earth remains relatively constant despite the production of new crust, a mechanism must exist by which crust is also destroyed.
+The destruction of oceanic crust occurs at subduction zones where oceanic crust is forced under either continental crust or oceanic crust.
+Today, the Atlantic basin is actively spreading at the Mid-Atlantic Ridge.
+Only a small portion of the oceanic;
+What is the answer of the following question:
+where does new crust come from in sea floor spreading
+PSR Prediction: the upper mantle
+BGM Context
+Given the passage titles and contexts below:
+Title: Seafloor spreading;
+Context: Seafloor spreading Seafloor spreading is a process that occurs at mid-ocean ridges, where new oceanic crust is formed through volcanic activity and then gradually moves away from the ridge.
+Earlier theories (e.g.
+by Alfred Wegener and Alexander du Toit) of continental drift postulated that continents "ploughed" through the sea.
+The idea that the seafloor itself moves (and also carries the continents with it) as it expands from a central axis was proposed by Harry Hess from Princeton University in the 1960s.
+The theory is well accepted now, and the phenomenon is known to be caused by convection currents in the;
+What is the answer of the following question:
+where does new crust come from in sea floor spreading
+BGM Prediction: volcanic activity
+Question II
+who has the most subscribers in faze clan
+Answer
+“FaZe Rug”
+GTR Context
+Given the passage titles and contexts below:
+Title: FaZe Clan;
+Context: "FaZe 2.0" by FaZe members and fans.
+The FaZe "CS:GO" team went on to become one of the most successful rosters for the 2017/2018 seasons.
+FaZe Clan is the most popular esports organization in the world, based on the organization’s social media following.
+As of November 28, 2018, FaZe Clan and its members together have 82 million YouTube subscribers, 11.2 billion YouTube views, 11.3 million Twitch followers, 130 million Twitch views, 43.1 million Twitter followers, 45.8 million Instagram followers, 2.8 million Facebook likes and followers.
+FaZe Clan has made \$6,148,290.91 from esport tournament prize pools alone.
+FaZe Clan started on;
+Title: FaZe Clan;
+Context: February 18, 2018.
+FaZe Clan FaZe Clan (formerly FaZe Sniping) is an American esports and entertainment organization that competes in various video game tournaments.
+The organization was founded as a gaming clan on YouTube by players known as Housecat, ClipZ, and Resistance in 2010, who all created "trickshot" videos for the video game "".
+In 2012, with the release of "", the organization decided to expand into competitive play.
+In 2016, a new era for FaZe began when the organization bought a "" professional team.
+This moment marked the beginning of FaZe Clan expanding into various esports.
+This movement is;
+Title: FaZe Clan;
+Context: FaZe Clan FaZe Clan (formerly FaZe Sniping) is an American esports and entertainment organization that competes in various video game tournaments.
+The organization was founded as a gaming clan on YouTube by players known as Housecat, ClipZ, and Resistance in 2010, who all created "trickshot" videos for the video game "".
+In 2012, with the release of "", the organization decided to expand into competitive play.
+In 2016, a new era for FaZe began when the organization bought a "" professional team.
+This moment marked the beginning of FaZe Clan expanding into various esports.
+This movement is referred to as;
+Title: FaZe Clan;
+Context: Duty" community has been frustrated with the way FaZe Clan has been/not been involved in the game in which it was founded.
+Most of the other FaZe members have changed their video topics and mostly do vlog videos.
+This sparked FaZe Clan to organize the highly successful FaZe Bootcamp on the release of "".
+FaZe trickshotters known as Kitty, Dirty, Bloo, GwidT and Replays contributed to a week dedicated to creating "Call of Duty" content.
+FaZe’s first roster—consisting of players named Heist, Folsom, Secretly, and Sham— was created to compete at the 2013 MLG Winter Championship.
+Their next roster—consisting of;
+Title: F Is for Family;
+Context: F Is for Family F is for Family is an American adult animated sitcom created by Bill Burr and Michael Price and produced by Gaumont International Television and Vince Vaughn’s Wild West Television.
+The show premiered on December 18, 2015, to generally favorable reviews.
+Season 2 premiered May 30, 2017.
+On June 28, 2017, the show was renewed for a third season.
+On July 1, 2018, Burr confirmed season three.
+On November 30, the third season was released.
+The series was announced in October 2014 as part of a partnership between Netflix, Gaumont International Television, and Wild West Television.;
+What is the answer of the following question:
+who has the most subscribers in faze clan
+GTR Prediction: Tfue
+PSR Context
+Given the passage titles and contexts below:
+Title: FaZe Clan;
+Context: Duty" community has been frustrated with the way FaZe Clan has been/not been involved in the game in which it was founded.
+Most of the other FaZe members have changed their video topics and mostly do vlog videos.
+This sparked FaZe Clan to organize the highly successful FaZe Bootcamp on the release of "".
+FaZe trickshotters known as Kitty, Dirty, Bloo, GwidT and Replays contributed to a week dedicated to creating "Call of Duty" content.
+FaZe’s first roster—consisting of players named Heist, Folsom, Secretly, and Sham— was created to compete at the 2013 MLG Winter Championship.
+Their next roster—consisting of;
+Title: FaZe Clan;
+Context: "FaZe 2.0" by FaZe members and fans.
+The FaZe "CS:GO" team went on to become one of the most successful rosters for the 2017/2018 seasons.
+FaZe Clan is the most popular esports organization in the world, based on the organization’s social media following.
+As of November 28, 2018, FaZe Clan and its members together have 82 million YouTube subscribers, 11.2 billion YouTube views, 11.3 million Twitch followers, 130 million Twitch views, 43.1 million Twitter followers, 45.8 million Instagram followers, 2.8 million Facebook likes and followers.
+FaZe Clan has made \$6,148,290.91 from esport tournament prize pools alone.
+FaZe Clan started on;
+Title: FaZe Clan;
+Context: FaZe Clan FaZe Clan (formerly FaZe Sniping) is an American esports and entertainment organization that competes in various video game tournaments.
+The organization was founded as a gaming clan on YouTube by players known as Housecat, ClipZ, and Resistance in 2010, who all created "trickshot" videos for the video game "".
+In 2012, with the release of "", the organization decided to expand into competitive play.
+In 2016, a new era for FaZe began when the organization bought a "" professional team.
+This moment marked the beginning of FaZe Clan expanding into various esports.
+This movement is referred to as;
+Title: F Is for Family;
+Context: F Is for Family F is for Family is an American adult animated sitcom created by Bill Burr and Michael Price and produced by Gaumont International Television and Vince Vaughn’s Wild West Television.
+The show premiered on December 18, 2015, to generally favorable reviews.
+Season 2 premiered May 30, 2017.
+On June 28, 2017, the show was renewed for a third season.
+On July 1, 2018, Burr confirmed season three.
+On November 30, the third season was released.
+The series was announced in October 2014 as part of a partnership between Netflix, Gaumont International Television, and Wild West Television.;
+Title: FaZe Clan;
+Context: February 18, 2018.
+FaZe Clan FaZe Clan (formerly FaZe Sniping) is an American esports and entertainment organization that competes in various video game tournaments.
+The organization was founded as a gaming clan on YouTube by players known as Housecat, ClipZ, and Resistance in 2010, who all created "trickshot" videos for the video game "".
+In 2012, with the release of "", the organization decided to expand into competitive play.
+In 2016, a new era for FaZe began when the organization bought a "" professional team.
+This moment marked the beginning of FaZe Clan expanding into various esports.
+This movement is;
+What is the answer of the following question:
+who has the most subscribers in faze clan
+PSR Prediction: Tfue
+BGM Context
+What is the answer of the following question:
+who has the most subscribers in faze clan
+BGM Prediction: FaZe Rug